OpenStudy (anonymous):

please help! Find the equation of a line that is perpendicular to the line y=1/3x+8, (6, 0) and contains the point (6, 0)

OpenStudy (texaschic101):

we need to find the slope in the equation given because a perpendicular line will have a negative reciprocal slope. So...do you know the slope of y = 1/3x + 8 ?

OpenStudy (texaschic101):

In y = mx + b form, the slope will be in the m position... y = mx + b y = (1/3)x + 8 what number is in the m position ?

OpenStudy (texaschic101):

when your ready to participate, let me know, and I will continue :)